=== 2019/1/17(Thur.) ===

=== Weekend ===
 * Nakano, Enomoto: Some work for Y arm alignment.

==== Transition from morning work (*:responsible person) ====

 * We have maintenance today.

==== Report of work (*:responsible person) ====

 * IOO: MC was recovered. Finesse was almost the same as before or slightly.
 * VIS: PR2 was almost recovered.
 * BS is almost ready.
 * AOS: WAB is being installed, needs alignment or something today.
 * SR: Measured SR hight was a bit lower than expected.

 * CAL: Y end PCal noise was improved drastically by 130dB as the same way of X end. We needs some small modifications for model or something.
 * DGS: Maintenance for RFM did not succeed to connect all center and both two ends. Needs investigation today.
 * DGS: X end rack maintenance will finish this Friday or beginning of next week.
 * VAC: Aluminum foil was removed at Y front. Only simple cover is placed to SR direction.

==== Work plan (*:responsible person) ====

 * COMM: Alignment will start this afternoon from FI. We need some iteration of alignment and adjustment of OSEM etc. for PRs and BS.